OpenStudy (anonymous):

The value of y varies directly with x. If y=80 when x is 16, what is x when y is 500? WILL GIVE MEDAl

Miracrown (miracrown):

If we know that y varies directly with x, it has to be a function like y=k*x and we need to find k. We're told that y=80 when x=16 if we plug those into our equation we get 80=16*k and we can solve that equation to find k What do you get?

OpenStudy (anonymous):

1/5

Miracrown (miracrown):

So you're telling me that, 16 x 1/5 = 80?

OpenStudy (anonymous):

no im just confused

Miracrown (miracrown):

16 multiplied by a number = 80 What is that ''number''?

OpenStudy (anonymous):

5

Miracrown (miracrown):

That's right! Good job! k=5 indeed so our function is y=5x

Miracrown (miracrown):

And if we want to know the value of x when y=500 We need to solve this equation: 500 = 5x So what's our answer?

OpenStudy (anonymous):

x=100

OpenStudy (anonymous):

Thanks

Miracrown (miracrown):

You're right x = 100 , good work! And you're most welcome :)